# References.bib

### v7

@@ 1.1 Contents
The contents of this study are divided into 5 parts, of which this paper is the first one. Each part contains several sections, and they are further subdivided into subsections. For example, you are currently reading Part 1, Section 1, Subsection 1.1.
Part 1 contains four main sections. Sections 3 and 5 are about foundational functions (F.F.) while Sections 4 and 6 are about redefining the Roman factorial recursively and non-recursively [^1] The term ”recursively”, as well as the concept of ”foundational functions”, will be clarified shortly. Any unfamiliar mathematical terminology introduced in this section or later will be explained in time.. This paper focuses on introducing the F.F. step by step, in a way that builds intuition and insight.
These F.F. are used for expressing the Roman factorial definition in a single formula (instead of a 2-part piece-wise definition) and in constructing a universal, non-recursive formula across all integers. The procedure of incorporating F.F. into expressions of the Roman factorial is referred to as the generalization process, and will be followed in Sections 4 and 6 as well as future papers.
@@
For those interested in diving deeper into specific mathematical concepts, an addendum is provided in Section 9. It has basic information about a variety of things and is intended to fill any gaps not explicitly elaborated upon in the main text. Lastly, Section 8 lists all references used in the formulation of this paper along with other details and information.
@@ 2.1 Factorial
Eq. 2.1.3 is called the recursive definition of the factorial, because in order to calculate $n!$ we need to know $(n-1)!\,$, starting with $0!=1$. Using this relationship we can find the next larger factorial for a given factorial. For example, if we know what $8!$ is, we can easily find $9!$ by evaluating for $n=9$: $9!=9\cdot 8!=362880$.
@@
If we substitute $n$ with $n+1$ in Eq. 2.1.3 and solve for $n!\,$, we get a similar equation [^6] In this paper, the symbol $\mathbb{Z}^{+}$ refers to natural numbers or positive integers $(1,2,3,\dots)$ while $\mathbb{Z}^{+}_{0}$ means non-negative integers $(0,1,2,3,\dots)$. More information about number sets in Addendum 9.1.:
$n!=\cfrac{(n+1)!}{n+1}\>,\quad n\in\mathbb{Z}^{+}_{0}\,.$ (Eq. 2.1.4)
@@
The factorial is depicted below in Fig. 2.1.1. Each point corresponds to integer factorials, while the Gamma function is used to connect these discrete points smoothly. This continuation will be presented in later parts of this study: its role in Fig. 2.1.1 is solely to be visually pleasing.
@@
When trying to find factorials of negative numbers, for example $(-1)!\,$, we come across a problem. We initially notice that there is no intuitive definition of the factorial for negative integers, so we need a new definition entirely or at least an expansion of the recursive relationship as iterated in Eq. 2.1.4.
Let’s approach this issue by attempting to find $(-1)!$ recursively. Using the recurrence relationship in Eq. 2.1.4 for $n=-1$, we observe that we have to divide by 0:
$(-1)!=\cfrac{(-1+1)!}{-1+1}=\cfrac{0!}{0}=\cfrac{1}{0}\>,$
@@
which is impossible since there is no value of $(-1)!$ that satisfies this equation (no number multiplied by 0 gives 1).
Furthermore, if we extend this process, $(-2)!$ is undefined and so are all negative integer factorials. If we attempt to evaluate them recursively, we come across other undefined values:
@@
$(-3)!$ $=\cfrac{(-3+1)!}{-3+1}=\cfrac{(-2)!}{-2}\enspace\cdots$
@@ 2.2 Roman factorial
The Roman factorial [roman] [romanfact] is an extension of the factorial function to negative integers [^7] The symbols [roman] and [romanfact] are references to citations, listed in Section 8.. It is named after Steven Roman, who used it in his work about umbral calculus. Roman factorial can be expressed in the following closed form:
@@
$n$7 -6 -5 -4 -3 -2 -1
$\lfloor n\rceil!$ $\nicefrac{{1}}{{720}}$$\nicefrac{{1}}{{120}}$ $\nicefrac{{1}}{{24}}$$\nicefrac{{1}}{{6}}$ $\nicefrac{{1}}{{2}}$1 1
@@
Negative Roman factorials exhibit a pattern akin to the reciprocals of positive factorials, with alternating signs. Notably, there is an offset: 5$\lfloor$-5$\rceil$!⌊ - 5 ⌉ ! is not the reciprocal of $5!$, but rather of $4!$. However, this offset does not affect any properties of the factorial.
@@
This relationship is defined for any integer $n$ except -1, which is expressed mathematically by the set [^8] More about number sets on Addendum 9.1. $\mathbb{Z}^{-}\setminus\{-1\}$. Eq. 2.2.2 does not define the Roman factorial recursively, it is just an identity valid for all integers except $-1$.
@@
Let’s confirm Eq. 2.2.3’s applicability to negative integers. Using 2$\lfloor$-2$\rceil$!⌊ - 2 ⌉ ! as a example, we can see that there is no issue in the calculation:
@@
The value of $(-2)!$ would normally be found using the recursive relationships of the normal factorial, if only $(-1)!$ was properly defined. 1$\lfloor$-1$\rceil$!⌊ - 1 ⌉ ! has a well-defined value only in the context of the Roman factorial, which provides values for all negative integers while being recursively consistent. Additionally, the division by $n+1$ is the reason for the alternating signs of negative integer factorials, since that divisor is negative when $n<-1$.
Certainly, there remains an anomaly in Eq. 2.2.2 particularly at $n=-1$, yet the piece-wise Roman factorial offers a solution: $\rceil!=1$ 1. This specific value doesn’t resolve the division by 0; indeed, no number can. Instead, the recursive relationship fails to provide a value for $(-1)!$ or 1$\lfloor$-1$\rceil$!⌊ - 1 ⌉ !, so we address this limitation by asserting that the relationship is undefined at this anomaly (assigning $(-1)!$ the value of 1 arbitrarily).
@@
The process of reaching this result will be explained thoroughly in Section 4 and it serves as an example of the generalization process used in this paper. This procedure seeks to rewrite various closed-form definitions as singular expressions by employing a set of foundational functions which have simple and analytic forms.
Sections 4 and 6 aim to establish two universal definitions for integer Roman factorials. One approach will result in a recursive definition and the other will result in a direct one, both applicable to all integers.
@@
To establish these definitions, we’ll introduce a set of foundational functions (F.F.) which will be explored in Sections 3 and 5 and will be used in the rest of the paper.
@@ 3.1 Introduction
Additionally, these output values will be expressed concisely as an "output pattern". Tbl. 3.1.1 can be represented by the output pattern $[0,1,0,-1,0]$. F.F. in this paper will be relevant for a few values of $n$, so the output tables will have 3 or so columns.
@@
This function can be expressed by the output pattern $[0,\nicefrac{{1}}{{e}},1,\nicefrac{{1}}{{e}},0]$. Most commonly though, output tables will have less columns. Using the previous example, we can condense Tbl. 3.1.2 like so:
$x$$\infty$ 0 +$\infty$+ ∞
@@
In this output table, the values for $\infty$- ∞ and +$\infty$+ ∞ are determined by taking the limit of $e^{-x^{2}}$ as $x$ approaches $\pm$ infinity. However, for the functions presented in this section, there is no need to take limits because they will output a single value across their entire domain. The example shown above serves only to illustrate the concept of output tables.
@@
In the rest of Section 3, as well as in Section 5, we will define a set of functions denoted by (F.F.). They will be helpful in Section 4 where we will condense Eq. 2.2.1 and Eq. 2.2.3 into a single relationship each, and in Section 6 in which we will build two universal product definitions for the Roman factorial. Foundational functions will be necessary for intuitively comprehending the mathematical definitions presented in this paper.
@@ 3.2 The function $\delta(n)$
$n$ldots -2 -1 0 1 2 $\ldots$
$\delta(n)$ldots -1.5 -0.5 0.5 1.5 2.5 $\ldots$
@@
Here, the symbols $+$ or -- denote any positive or negative number, respectively. The function $\delta(n)$ aims to yield outputs with specific signs, the exact value of which is not relevant. Additionally, $\varepsilon$ cannot take the value of 0 or 1, because that would mean either $\delta(0)=0$ or $\delta(-1)=0$, and this is not desired.
@@
$n$$\infty$ $\ldots$ -2 -1 0 1 2 $\ldots$ $\infty$
$\delta(n)$- $\ldots$ -- -- + + + $\ldots$ +
@@
Hence, $\delta(n)$ exhibits the output pattern [,+,+][-,+,+][ - , + , + ]. The parameter $\varepsilon$ can assume any value between 0 and 1 while adhering to the specified behavior.
@@ 3.3 The function $\theta(n)$
The second F.F., denoted by $\theta(n)$, is simple. Its primary objective is to transform the positive and negative outputs of $\delta(n)$ into $+1$ or $-1$, respectively.
@@
$\theta(n)=\frac{\delta(n)}{\left|\delta(n)\right|}\enspace,\quad n\in\mathbb{R}\,.$ (Eq. 3.3.2)
@@
While resembling the $sign$ function’s functionality, $\theta(n)$ is specifically chosen due to a critical distinction: $sign(0)=0$, whereas $\theta(0)=1$. This difference is important, and it is the rationale behind the definition of $\delta(n)$. A function similar to the $sign$ function, yet distinct from outputting zero, is necessary. The function $\theta(n)$ has the output pattern $[-1,1,1]$ and is crucial for defining F.F. as well as building a universal Roman factorial in Section 4.
@@ 3.4 The function $\xi(n)$
Let’s construct $\xi(n)$ using $\theta(n)$ step by step. To begin with, we note that the output pattern of $\theta(n)$ is $[-1,1,1]$. Adding 1 to $\theta(n)$ essentially adds to each output, obtaining $[0,2,2]$. Dividing $\big{(}1+\theta(n)\big{)}$ by 2 yields the desired pattern $[0,1,1]$. Thus, we arrive at the universal definition of $\xi(n)$:
@@
This function has similarities to true/false states in Boolean logic, which is commonly used in programming. Specifically, the expressions $\xi(n\geq 0)=1$ and $\xi(n<0)=0$ represent the true/false statements of the condition $n\geq 0$. Furthermore, it’s noteworthy that the most subsequent F.F. can also be expressed in Boolean algebra [^11] More about Boolean algebra in Addendum 9.5..
@@ 3.5 The function $\xi^{\prime}(n)$
In Boolean algebra, $\xi^{\prime}(n)$ yields the true/false states of the condition $n<0\,$, thereby exhibiting an output pattern of $[1,0,0]$.
@@ 3.6 The function $\eta(n)$
This almost works, however the signs are opposite of what we specified in Eq. 3.6.1. This can be easily solved by raising $(-1)$ to the power $(n-1)$:
@@
Let’s now address the case of positive inputs. Since we raise the base $(-1)$ to some power, if the base becomes $(+1)$ for $n>0$, then the result will always be 1 regardless of the power. That is true because $1^{n}=1,\enspace n\in\mathbb{R}\,$. To ensure that behavior we can use $\theta(n)$ for the base:
@@
Also, by adjusting the exponent to $(\lceil n\rceil-1)$, where $\lceil n\rceil$ is the ceiling function (rounding up), it is guaranteed that $(-1)$ is not raised to some fractional power for non-integer $n$. This initiative originates from the aim to define F.F. for all real numbers, not just integers.
@@
In the introduction to the Roman factorial, we observed a term similar to $\eta(n)$ in the piece-wise definition of Eq. 2.2.1. However, instead of having the power $(n-1)$, it has $(-n-1)$. Since we aim for $\eta(n)$ to be part of the universal Roman factorial definition, these terms should match exactly.
Fortunately, we can simply substitute $n$ with $-n$ directly within the power term of $\eta(n)$. This is because it doesn’t matter if $(-1)$ is raised to a positive or negative power:
@@ 3.7 Synopsis
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
@@
You can refer to these equations for a reminder of the output patterns when these functions are utilized in Section 4. Additional F.F. will be introduced in Section 5, to be employed in Section 6.
@@ 4.1 Introduction
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
@@
Despite the rewriting, the resulting relationships will still be based on a recursive factorial. Later, in Section 6, we will develop a new factorial definition from scratch.

### v8

# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

@@ 1.1 Contents
The contents of this study are divided into 5 parts, of which this paper is the first one. Each part contains several sections, and they are further subdivided into subsections. For example, you are currently reading Part 1, Section˜1, Subsection˜1.1.
Part 1 contains four main sections. Sections˜3 and 5 are about foundational functions (F.F.) while Sections˜4 and 6 are about redefining the Roman factorial recursively and non-recursively [^1] The term ”recursively”, as well as the concept of ”foundational functions”, will be clarified shortly. Any unfamiliar mathematical terminology introduced in this section or later will be explained in time.. This paper focuses on introducing the F.F. step by step, in a way that builds intuition and insight.
These F.F. are used for expressing the Roman factorial definition in a single formula (instead of a 2-part piece-wise definition) and in constructing a universal, non-recursive formula across all integers. The procedure of incorporating F.F. into expressions of the Roman factorial is referred to as the generalization process, and will be followed in Sections˜4 and 6 as well as future papers.
@@
For those interested in diving deeper into specific mathematical concepts, an addendum is provided in Section˜9. It has basic information about a variety of things and is intended to fill any gaps not explicitly elaborated upon in the main text. Lastly, Section˜8 lists all references used in the formulation of this paper along with other details and information.
@@ 2.1 Factorial
Eq. 2.1.3 is called the recursive definition of the factorial, because in order to calculate $n!$ we need to know (n1)!(n-1)!\,, starting with $0!=1$. Using this relationship we can find the next larger factorial for a given factorial. For example, if we know what $8!$ is, we can easily find $9!$ by evaluating for $n=9$: $9!=9\cdot 8!=362880$.
@@
If we substitute $n$ with $n+1$ in Eq. 2.1.3 and solve for n!n!\,, we get a similar equation [^6] In this paper, the symbol $\mathbb{Z}^{+}$ refers to natural numbers or positive integers $(1,2,3,\dots)$ while $\mathbb{Z}^{+}_{0}$ means non-negative integers $(0,1,2,3,\dots)\,$. More information about number sets in Addendum 9.1.:
$n!=\cfrac{(n+1)!}{n+1}\;,\quad n\in\mathbb{Z}^{+}_{0}\,$. (Eq. 2.1.4)
@@
The factorial is depicted below in ˜Fig. 2.1.1. Each point corresponds to integer factorials, while the Gamma function is used to connect these discrete points smoothly. This continuation will be presented in later parts of this study: its role in ˜Fig. 2.1.1 is solely to be visually pleasing.
@@
When trying to find factorials of negative numbers, for example (1)!(-1)!\,, we come across a problem. We initially notice that there is no intuitive definition of the factorial for negative integers, so we need a new definition entirely or at least an expansion of the recursive relationship as iterated in Eq. 2.1.4.
Let’s approach this issue by attempting to find (1)!(-1)! recursively. Using the recurrence relationship in Eq. 2.1.4 for n=1n=-1, we observe that we have to divide by 0:
$(-1)!=\cfrac{(-1+1)!}{-1+1}=\cfrac{0!}{0}=\cfrac{1}{0}\;,$
@@
which is impossible since there is no value of (1)!(-1)! that satisfies this equation (no number multiplied by 0 gives 1).
Furthermore, if we extend this process, (2)!(-2)! is undefined and so are all negative integer factorials. If we attempt to evaluate them recursively, we come across other undefined values:
@@
$(-3)!$ $=\cfrac{(-3+1)!}{-3+1}=\cfrac{(-2)!}{-2}\>\cdots$
@@ 2.2 Roman factorial
The Roman factorial [1] [2] is an extension of the factorial function to negative integers [^7] The symbols [1] and [2] are references to citations, listed in Section 8.. It is named after Steven Roman, who used it in his work about umbral calculus. Roman factorial can be expressed in the following closed form:
@@
$n$ -7 -6 -5 -4 -3 -2 -1
$\lfloor n\rceil!$ $\nicefrac{{1}}{{720}}$ -$\nicefrac{{1}}{{120}}$ $\nicefrac{{1}}{{24}}$ -$\nicefrac{{1}}{{6}}$ $\nicefrac{{1}}{{2}}$ -1 1
@@
Negative Roman factorials exhibit a pattern akin to the reciprocals of positive factorials, with alternating signs. Notably, there is an offset: $\lfloor-5\rceil!$ is not the reciprocal of $5!$, but rather of $4!$. However, this offset does not affect any properties of the factorial.
@@
This relationship is defined for any integer $n$ except -1, which is expressed mathematically by the set [^8] More about number sets on Addendum 9.1. $\mathbb{Z}^{-}\setminus\{-1\}$. Eq. 2.2.2 does not define the Roman factorial recursively, it is just an identity valid for all integers except 1-1.
@@
Let’s confirm Eq. 2.2.3’s applicability to negative integers. Using $\lfloor-2\rceil!$ as a example, we can see that there is no issue in the calculation:
@@
The value of (2)!(-2)! would normally be found using the recursive relationships of the normal factorial, if only (1)!(-1)! was properly defined. $\lfloor-1\rceil!$ has a well-defined value only in the context of the Roman factorial, which provides values for all negative integers while being recursively consistent. Additionally, the division by $n+1$ is the reason for the alternating signs of negative integer factorials, since that divisor is negative when n<1n<-1.
Certainly, there remains an anomaly in Eq. 2.2.2 particularly at n=1n=-1, yet the piece-wise Roman factorial offers a solution: $\lfloor-1\rceil!=1$. This specific value doesn’t resolve the division by 0; indeed, no number can. Instead, the recursive relationship fails to provide a value for (1)!(-1)! or $\lfloor-1\rceil!$, so we address this limitation by asserting that the relationship is undefined at this anomaly (assigning (1)!(-1)! the value of 1 arbitrarily).
@@
The process of reaching this result will be explained thoroughly in Section˜4 and it serves as an example of the generalization process used in this paper. This procedure seeks to rewrite various closed-form definitions as singular expressions by employing a set of foundational functions which have simple and analytic forms.
Sections˜4 and 6 aim to establish two universal definitions for integer Roman factorials. One approach will result in a recursive definition and the other will result in a direct one, both applicable to all integers.
@@
To establish these definitions, we’ll introduce a set of foundational functions (F.F.) which will be explored in Sections˜3 and 5 and will be used in the rest of the paper.
@@ 3.1 Introduction
Additionally, these output values will be expressed concisely as an "output pattern". Tbl. 3.1.1 can be represented by the output pattern [0,1,0,1,0][0,1,0,-1,0]. F.F. in this paper will be relevant for a few values of $n$, so the output tables will have 3 or so columns.
@@
This function can be expressed by the output pattern $0,\nicefrac{{1}}{{e}},1,\nicefrac{{1}}{{e}},0]$. Most commonly though, output tables will have less columns. Using the previous example, we can condense Tbl. 3.1.2 like so:
$x$ -$\infty$ 0 $\infty$
@@
In this output table, the values for -$\infty$ and $\infty$ are determined by taking the limit of $e^{-x^{2}}$ as $x$ approaches $\pm$ infinity. However, for the functions presented in this section, there is no need to take limits because they will output a single value across their entire domain. The example shown above serves only to illustrate the concept of output tables.
@@
In the rest of Section˜3, as well as in Section˜5, we will define a set of functions denoted by (F.F.). They will be helpful in Section˜4 where we will condense Eq. 2.2.1 and Eq. 2.2.3 into a single relationship each, and in Section˜6 in which we will build two universal product definitions for the Roman factorial. Foundational functions will be necessary for intuitively comprehending the mathematical definitions presented in this paper.
@@ 3.2 The function $\delta(n)$
$n$ $\ldots$ -2 -1 0 1 2 $\ldots$
$\delta(n)$ $\ldots$ -1.5 -0.5 0.5 1.5 2.5 $\ldots$
@@
Here, the symbols ++ or - denote any positive or negative number, respectively. The function $\delta(n)$ aims to yield outputs with specific signs, the exact value of which is not relevant. Additionally, $\varepsilon$ cannot take the value of 0 or 1, because that would mean either $\delta(0)=0$ or $\delta(-1)=0$, and this is not desired.
@@
$n$ -$\infty$ $\ldots$ -2 -1 0 1 2 $\ldots$ $\infty$
$\delta(n)$ - $\ldots$ - - + + + $\ldots$ +
@@
Hence, $\delta(n)$ exhibits the output pattern [,+,+][-,+,+]. The parameter $\varepsilon$ can assume any value between 0 and 1 while adhering to the specified behavior.
@@ 3.3 The function $\theta(n)$
The second F.F., denoted by $\theta(n)$, is simple. Its primary objective is to transform the positive and negative outputs of $\delta(n)$ into +1+1 or 1-1, respectively.
@@
$\theta(n)=\frac{\delta(n)}{\left|\delta(n)\right|}\>,\quad n\in\mathbb{R}\,.$ (Eq. 3.3.2)
@@
While resembling the signsign function’s functionality, $\theta(n)$ is specifically chosen due to a critical distinction: $sign(0)=0$, whereas $\theta(0)=1$. This difference is important, and it is the rationale behind the definition of $\delta(n)$. A function similar to the signsign function, yet distinct from outputting zero, is necessary. The function $\theta(n)$ has the output pattern [1,1,1][-1,1,1] and is crucial for defining F.F. as well as building a universal Roman factorial in Section˜4.
@@ 3.4 The function $\xi(n)$
Let’s construct $\xi(n)$ using $\theta(n)$ step by step. To begin with, we note that the output pattern of $\theta(n)$ is [1,1,1][-1,1,1]. Adding 1 to $\theta(n)$ essentially adds to each output, obtaining $[0,2,2]$. Dividing $\big{(}1+\theta(n)\big{)}$ by 2 yields the desired pattern $[0,1,1]$. Thus, we arrive at the universal definition of $\xi(n)$:
@@
This function has similarities to true/false states in Boolean logic, which is commonly used in programming. Specifically, the states $\xi(n\geq 0)=1$ and $\xi(n<0)=0$ represent the true/false statements of the condition $n\geq 0$. Furthermore, it’s noteworthy that the most subsequent F.F. can also be expressed in Boolean algebra [^11] More about Boolean algebra in Addendum 9.5..
@@ 3.5 The function $\xi^{\prime}(n)$
In Boolean algebra, $\xi^{\prime}(n)$ yields the true/false states of the condition n<0n<0\,, thereby exhibiting an output pattern of $[1,0,0]$.
@@ 3.6 The function $\eta(n)$
This almost works, however the signs are opposite of what we specified in Eq. 3.6.1. This can be easily solved by raising (1)(-1) to the power (n1)(n-1):
@@
Let’s now address the case of positive inputs. Since we raise the base (1)(-1) to some power, if the base becomes $(+1)$ for $n>0$, then the result will always be 1 regardless of the power. That is true because $1^{n}=1,\enspace n\in\mathbb{R}\,$. To ensure that behavior we can use $\theta(n)$ for the base:
@@
Also, by adjusting the exponent to $(\lceil n\rceil-1)$, where $\lceil n\rceil$ is the ceiling function (rounding up), it is guaranteed that (1)(-1) is not raised to some fractional power for non-integer $n$. This initiative originates from the aim to define F.F. for all real numbers, not just integers.
@@
In the introduction to the Roman factorial, we observed a term similar to $\eta(n)$ in the piece-wise definition of Eq. 2.2.1. However, instead of having the power (n1)(n-1), it has (n1)(-n-1). Since we aim for $\eta(n)$ to be part of the universal Roman factorial definition, these terms should match exactly.
Fortunately, we can simply substitute $n$ with n-n directly within the power term of $\eta(n)$. This is because it doesn’t matter if (1)(-1) is raised to a positive or negative power:
@@ 3.7 Synopsis
$\delta(n)$ $=\lfloor n\rfloor+0.5$ $[-,+,+]$ (Eq. 3.2.3)
@@
You can refer to these equations for a reminder of the output patterns when these functions are utilized in Section˜4. Additional F.F. will be introduced in Section˜5, to be employed in Section˜6.
@@ 4.1 Introduction
[…]
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
@@
Despite the rewriting, the resulting relationships will still be based on a recursive factorial. Later, in Section˜6, we will develop a new factorial definition from scratch.